QUESTION 4 PARTNERSHIPS [20 MARKS REQUIRED Use the information provided below to prepare the Statement of Changes in Equity for the year ended 30 June 2014. Use the following format STATEMENT OF CHANGES IN EQUITY FOR THE YEAR ENDED 30 JUNE 2014 Fuji Film Total Capital Accounts Balance at 30 June 2013 ges in capita Balance at 30 June 2014 Total Current Accounts Fuji Film ance at 30 June 2013 Net profit for the year Interest on Capital Bonus Profit Share Balance at 30 June 2014 INFORMATION Extract from the ledger of Fuji-Film Traders on 30 June 2014 Capital: Fuj 350 000 tal: Film 100 Current ac: Fu 50 000 talc: Film 20 000 Drawings: Fuji 90000 Drawings: Film 40 000
The following must be taken into account: (a) On 30 June 2014 the net profit for the year amounted to R520 000 (b) The partnership agreement makes provision for the following: Interest on capital must be provided at 12% per annum on the balances in the capital accounts for the first 6 months of the year. The nterest rate increases to 15% for the last 6 months of the year The partners are entited to the following annual salaries: Fuj R 25 000 Film R 20 000 Film is entitled to a special bonus equal to 5% of the net profit before any of the above * appropriations have been taken into account. Fuji decreased his capital contribution by R50 000 on 1 January 2014. This capital decrease has been recorded. Fuji and Film share the remaining profits or losses in the ratio of 3:2 respectively.
Answer #1

STATEMENT OF CHANGES IN EQUITY AS ON 30 JUNE 2014

CAPITAL ACCOUNTS

FUJI

FLIM

TOTAL

BALANCE AT 30JUN 2013

400000

100000

500000

CHANGES IN CAPITAL

50000

-

50000

BALANCE AT 30JUN 2014

350000

100000

450000

CURRENT ACCOUNTS

FUJI

FLIM

TOTAL

BALANCE AT 30JUN 2013

166400

(Dr. BALANCE)

153600

(Dr. BALANCE)

320000

(Dr. BALANCE)

NET PROFIT

520000

INTEREST ON CAPITAL

50250

13500

63750

SALARIES

25000

20000

45000

BONUS

-

26000

26000

PROFIT SHARE

231150

154100

385250

DRAWINGS

90000

40000

130000

BALANCE AT 30JUN 2014

50000

20000

70000


INTEREST ON CAPITAL

FIRST 6 MONTH 12% PAR ANNUAM

FUJI- CAPITAL BALANCE AS ON 31 DEC 2014 -400000*12%*6/12=24000

NEXT 6 MONTH 15% PER ANNUAM

FUJI- CAPITAL BALANCE AS ON 30JUN 2014-350000*15%*6/12=26250

TOTAL    50250

FLIM- CAPITAL BALANCE AS ON 31 DEC 2014 -100000*12%*6/12=6000

FLIM- CAPITAL BALANCE AS ON 30JUN 2014- 100000*15%*6/12=7500

TOTAL                                                                                                 13500

COMPUTATION OF PROFIT SHARE

NET PROFIT---                                                                                  520000

Less:- INTEREST ON CAPITAL- 63750

SALARIES- 45000

BONUS-   26000

DIVISIBLE PROFIT- 385,250

PROFIT SHARE OF FUJI= 385,250*3/5=                                    231,150

PROFIT SHARE OF FLIM= 385,250*2/5=                                   154,100